# Quantifying the Impact of the Dust Torque on the Migration of Low-mass Planets II: The Role of Pebble Accretion in Planet Growth within a Global Planet Formation Model

### v1

@@ 3.4 Planet Formation Tracks in Metal-rich Disks
This increased dust-to-gas ratio substantially affects the formation tracks of planets initiating their formation inside the water ice-line (specifically, those at 0.5, 1, and 2 au, as shown in the top panel of Figure 9). In these instances, when the dust torque is accounted for (black lines), all these planets migrate outward during the early stages of their development. Unlike the previous cases, these outward-migrating planets ultimately attain greater masses than they would have reached had the dust torque been ignored. Notably, the planet starting at 2 au crosses the water ice-line, sparking rapid formation through the accretion of pebbles with larger Stokes numbers (as highlighted in the bottom panel of Figure 10).
@@
Finally, for planets that are initially situated beyond the water ice-line (those positioned beyond 3 au in Figure 9), we once again observe that the dust torque has minimal impact. Its presence only slightly delays inward migration, resulting in planet formation tracks that remain largely unchanged.
Figure 8: Time evolution maps of the particle mass-weighted mean Stokes numbers (top) and the dust-to-gas mass ratio $\epsilon=\Sigma_{\text{d}}/\Sigma_{\text{g}}$ (bottom) for the case of a metallic disk with an initial dust-to-gas ratio of 0.02.
Figure 9: Top panel: same as Figure 2 but adopting an initial dust-to-gas ratio of 0.02. Bottom panel: time evolution of the core mass (solid lines) and envelope mass (dashed lines) for the planet initially located at 1 au.
@@ 4 The Impact of Pebble Accretion
The first attempt to incorporate pebble accretion in the context of pebble-driven torques was made by Regály (2020). In their multi-fluid model, pebble accretion was implemented by artificially reducing the pebble surface density, while maintaining a fixed accretion radius and treating the accretion efficiency as a variable parameter. This approach effectively sidestepped self-consistent modeling of pebble trajectories converging at the planet’s location. Despite this simplification, Regály (2020) demonstrated that for low Stokes numbers and low-mass planets, including pebble accretion can lead to a significantly strong positive dust torque. A similar finding was reported more recently by Chrenko et al. (2024), who investigated the effect of incorporating pebble accretion on the dust torque using a new hybrid method that allows pebble trajectories to converge to the planet’s radius (as shown in their Figures 2 and 3). Furthermore, through their Equations (16) to (18), Chrenko et al. (2024) provided a relatively straightforward recipe to fit their results.
@@
We now compare the planet formation tracks generated using the results from BLP18 and those from Chrenko et al. (2024), who used an analytical formula to fit their numerical results. In Figure 11, we present the planet formation tracks for the fiducial case, contrasting scenarios that do not consider the dust torque (red lines) with those that incorporate results from BLP18 (black solid lines) and the tracks obtained by implementing in our code the new recipes provided in Chrenko et al. (2024) (violet solid lines). For planets initially located beyond the water ice-line (i.e., beyond 3 au), Figure 11 illustrates that, although the overall formation tracks remain similar, the dust torque appears more effective at greater distances when employing the prescriptions of Chrenko et al. (2024). This results in a slightly more pronounced delay of rapid inward migration for those planets. In contrast, for planets that begin their formation inside the water ice-line, the effects of the dust torque become markedly more significant. This is because, according to the results in Chrenko et al. (2024) the positive magnitude of the dust torque is notably greater for low-mass planets and small Stokes numbers when pebble accretion is taken into account. Consequently, these planets migrate outward significantly during the early stages of formation. In addition, in this case all the inner planets cross the water-ice line and their formation is triggered by the accretion of ice-rich pebbles with larger Stokes numbers (as in the case of the metallic disk). Eventually, at some point along their formation pathways, the gas torque begins to dominate, leading the planets to migrate inward.
This outcome –that the dust torque is more efficient for inner planets– aligns with the estimates provided by Chrenko et al. (2024), who employed a uniform pebble flux (see their Figure 7).
@@
Figure 11: Planet formation tracks for the fiducial disk –as in Figure 2– including now the impact of pebble accretion onto the dust torque (violet lines).
@@ 5 Summary and Discussion
The studies by BLP18, Regály (2020), and, more recently by Chrenko et al. (2024), have demonstrated that the asymmetric distribution of pebbles surrounding a planet can exert a significant torque on it, even though solid materials make up only a small fraction of the protoplanetary disk’s mass. Building on these findings (as well as our previous work, Guilera et al., 2023), in this study we have quantified for the first time the influence of dust torque on the migration of planets growing through pebble accretion until either the disk dissipates or, conservatively, the planets reach a mass of $10~{}M_{\oplus}$. We accomplished this goal with PLANETALP, a comprehensive framework for modeling the formation history of planets that incorporates dust growth and evolution (Ronco et al., 2017; Guilera et al., 2017, 2019; Venturini et al., 2020b; Guilera et al., 2021).
We first employed a fiducial disk model with a moderate mass of $0.05M_{\oplus}$ and a low $\alpha$-viscosity parameter of $10^{-4}$. Previous works on dust growth and evolution suggest that the efficiency of planet formation via pebble accretion is highly sensitive to the level of turbulence in the disk (e.g., Venturini et al., 2020b, a; Drazkowska et al., 2021; Guilera et al., 2021). On observational grounds, ALMA data suggest low $\alpha$-viscosity values (e.g., Dullemond et al., 2018; Flaherty et al., 2020). Therefore, based on these results and the findings of Venturini et al. (2020a) and Guilera et al. (2021), we considered a low $\alpha$-viscosity value, which enhances planetary formation efficiency.
@@
Impact of Dust Torque in Massive and Metal Rich Disks – We also investigated the role of the dust torque in the migration of planets within disks characterized by higher masses and metallicities. In both scenarios, we found that the pebble Stokes numbers across the disk remained comparable to those in the fiducial case. This similarity arises from the fact that dust growth is limited by fragmentation in the inner disk and by radial drift in the outer regions, resulting in minimal differences between these models. Consequently, the outcomes for these scenarios qualitatively mirror those of the fiducial case. Although the dust torque plays a crucial role in the formation of planets initially located inside the water ice-line, it is less significant for those beginning their formation beyond this line. The consistency of this finding is primarily due to the increased dust-to-gas mass ratio inside the water ice-line and the corresponding decrease outside, as a consequence of dust radial drift.
Notably, the disk with an initial metallicity of $Z=0.02$ exhibits a higher dust-to-gas ratio, resulting in a more pronounced impact of dust torque. In this case, planets starting at 0.5 au and 1 au migrate outward until they reach approximately $2~{}M_{\oplus}$, while the planet initially at 2 au crosses the water ice-line. This change not only affects the final mass of the planet but also alters its composition. A similar trend was observed in Guilera et al. (2021) when updated prescriptions for thermal torque were implemented.
@@
For planets initiating their formation within the water ice line, the influence of dust torque becomes more pronounced compared to the findings from BLP18. This is primarily because Chrenko et al. (2024) demonstrated that, when considering pebble accretion, the dust torque tends to increase for low-mass planets with low Stokes numbers.
For planets initially positioned beyond the water ice line, we observe that considering the effects of pebble accretion on the dust torque has minimal influence on the formation tracks. This is even though the dust torque incorporating accretion is larger in magnitude than the one calculated using BLP18’s results. This occurs primarily due to the brief timescales for planet formation and the rapid decrease in the dust-to-gas mass ratio beyond the water-ice line. Additionally, we observed that the effect of dust torque incorporating accretion appears to be amplified at greater distances. Moreover, we note that the impact of the dust torque, when including accretion, seems to increase at larger distances from the star. This enhancement might stem from the different dependence of the dust torque on the normalized torque parameters, contrasting our approach (using the results from BLP18) with those from Chrenko et al. (2024, see their discussion in Sec. 3.5).
@@
We also note that the planets we modeled start forming at the beginning of the simulations, specifically at $t=0$ yr. As demonstrated by Drazkowska et al. (2021) and Guilera et al. (2021), if planetary embryos do not start forming in the early phases of disk evolution, pure pebble accretion becomes inefficient due to diminishing pebble flux over time caused by dust radial drift. More recently, Voelkel et al. (2022) incorporated a model for planetesimal formation through pebble flux and embryo formation from planetesimal growth in a framework similar to the one we use in this study. Their findings revealed that multiple generations of embryos can form prior to disk dissipation, particularly within the water ice-line, and that earlier generations primarily grow via pebble accretion, while later ones tend to grow through planetesimal accretion. We plan to incorporate these phenomena in future studies to examine further the impact of the dust torque on different generations of formed planets.
@@ 6 Conclusions
Planets initially inside the water ice-line – In this case, the dust torque represents a significant contribution to the total torque due to the increased dust-to-gas mass ratio resulting from dust drift from the outer disk. Moreover, at early stages, the dust torque can dominate the total torque acting on the planet, particularly in disks with initial metallicities slightly above the standard 1% ($\epsilon_{0}\gtrsim 0.02$). However, to achieve more accurate results, further detailed simulations extending to lower planet masses and Stokes numbers are necessary, especially when applying the recent results from Chrenko et al. (2024). For these planets, we identify three primary pathways:
@@
iii) when applying the prescriptions from Chrenko et al. (2024), all cases show that the planets achieve greater masses. In this scenario, the dust torque plays a more significant role in moving the planets outward, enabling them to cross the water-ice line and facilitating their formation through the accretion of larger pebbles, which also alters the planets’ composition by incorporating ice-rich pebbles.
@@
In conclusion, we assert that the dust torque should not be ignored in models of planet formation, especially in research centered on the development of terrestrial-like planets through pebble accretion. In this regard, the dust torque can significantly influence the overall torque experienced by planets undergoing pebble accretion, thereby impacting their migration patterns and, ultimately, their formation timescale, mass, and composition.
@@ 7 Acknowledgments
OMG and M3B are partially supported by PIP 2971 from CONICET (Argentina) and by PICT 2020-03316 from ANPCyT (Argentina). OMG and M3B also thank Juan Ignacio Rodriguez from IALP for the computation managing resources of the Grupo de Astrofísica Planetaria de La Plata. PBL acknowledges support from FONDECYT project 1231205. The research leading to this work was supported by the Independent Research Fund Denmark via grant ID 10.46540/3103-00205B attributed to MP.
@@ Appendix A Dust torque from the dust distribution
In this appendix, we now compute the dust torque from the dust distribution instead of using only the mass-weighted mean Stokes number of such distribution. Even though it is a more expensive computational approach, it is useful to analyze if there is a remarkable difference with respect to the most simple case of considering a representative Stokes number for the dust distribution. To do this, we compute now the dust torque at the location of the planet for the 200 dust species, which are logarithmically equally spaced between $\text{r}_{\text{d}}^{\text{min}}=1~{}\mu\text{m}$ ($\text{St}_{\text{min}}$) and $\text{r}_{\text{d}}^{\text{max}}$ ($\text{St}_{\text{max}}$) (see Guilera et al., 2020, for details). We note again that in our model – which is based on the works of Birnstiel et al. (2012) and Dra̧zkowska et al. (2016) – most of the mass remains in larger particles. We compute the dust-to-gas ratio corresponding to the i-species given by
@@
In Figure 12, we show a comparison of the planet formation tracks for the case where the dust torque is computed using the mass-weighted mean Stokes number of the dust distribution (the black lines) and the case where the dust torque is calculated for each dust species (the multi-species case represented by the blue lines). We also plot the case where the dust torque is not consider in the total torque over the planets (the red lines). For the planets initially located inside the water ice-line, the planet formation tracks for the cases when the dust torque is considered are identical. This is due to the fact that inside the water ice-line the maximum Stokes number $\text{St}^{\text{max}}$ of the dust distribution is always lower than 0.01 (see top panel of Figure 8). This also implies that the mass-weighted mean Stokes number is also lower than 0.01. As we are not extrapolating outside the values of the dust torque map computed in BLP18, for St lower than 0.01 we are using this value to compute the dust torque. Hence, in the multi-species approach the dust torque for each dust species is the same, and the total dust torque becomes the same that the dust torque computed using mass-weighted mean Stokes number. On the other hand, beyond the water ice-line the maximum Stokes number $\text{St}^{\text{max}}$ of the dust distribution reach values larger than 0.01. Thus, in the multi-species approach not all the species contribute with the same value of the dust torque (particles with larger St contribute with larger dust torques). For those planets initially located beyond the water ice-line, we can see that there is a slightly difference between the planet formation tracks when the dust torque is considered. The choice to compute the dust torque using the mass-weighted mean Stokes number of the dust distribution slightly overestimates the effect of the dust torque respect to the case of computing the total dust torque considering the full dust distribution. However, in all cases the dust torque delays the inward planet migration respect to the cases where the dust torque is not considered.
Figure 12: Comparison of the planet formation tracks for the case of an initial disk metallicity of 0.02. The red curves correspond to case where the total torque is given just by the gaseous torque (where we use the recipes from Jiménez & Masset, 2017). The black and blue lines correspond to the case where the dust torque is considered. The black lines represent the case where the dust torque is computed adopting the mass-weighted mean Stokes number of the dust distribution, while the blue ones represent the case where the dust torque is computed for the 200 species of such distribution.

### v2

@@ 3.4 Planet Formation Tracks in Metal-rich Disks
This increased dust-to-gas ratio substantially affects the formation tracks of planets initiating their formation inside the water ice-line (specifically, those at 0.5, 1, and 2 au, as shown in the top panel of Figure 9). In these instances, when the dust torque is accounted for (black lines), all these planets migrate outward during the early stages of their development. Unlike the previous cases, these outward-migrating planets ultimately attain larger masses than they would have reached had the dust torque been ignored. Notably, the planet starting at 2 au crosses the water ice-line, sparking rapid formation through the accretion of pebbles with larger Stokes numbers (as highlighted in the bottom panel of Figure 10).
@@
Finally, for planets that are initially situated beyond the water ice-line (those positioned beyond 3 au in Figure 9), we once again observe that the dust torque has minimal impact. The dust torque only has a minor effect on delaying inward migration, leading to planet formation paths that remain mostly the same.
Figure 8: Time evolution maps of the particle mass-weighted mean Stokes numbers (top) and the dust-to-gas mass ratio $\epsilon=\Sigma_{\text{d}}/\Sigma_{\text{g}}$ (bottom) for the case of a metallic disk with an initial dust-to-gas ratio $\epsilon=0.02$.
Figure 9: Top panel: same as Figure 2 but adopting an initial dust-to-gas ratio $\epsilon=0.02$. Bottom panel: time evolution of the core mass (solid lines) and envelope mass (dashed lines) for the planet initially located at 1 au.
@@ 4 The Impact of Pebble Accretion
The first attempt to incorporate pebble accretion in the context of pebble-driven torques was made by Regály (2020). In their multi-fluid model, pebble accretion was implemented by artificially reducing the pebble surface density, while maintaining a fixed accretion radius and treating the accretion efficiency as a variable parameter. This approach effectively sidestepped self-consistent modeling of pebble trajectories converging at the planet’s location. In spite of this simplification, Regály (2020) demonstrated that for low Stokes numbers and low-mass planets, including pebble accretion can lead to a significantly strong positive dust torque. A similar finding was reported more recently by Chrenko et al. (2024), who investigated the effect of incorporating pebble accretion on the dust torque using a new hybrid method that allows pebble trajectories to converge to the planet’s radius (as shown in their Figures 2 and 3). Furthermore, through their Equations (16) to (18), Chrenko et al. (2024) provided a relatively straightforward recipe to fit their results.
@@
It is interesting to compare the planet formation tracks generated using the results from BLP18 and those from Chrenko et al. (2024), who used an analytical formula to fit their numerical results. In Figure 11, we present the planet formation tracks for the fiducial case, contrasting scenarios that do not consider the dust torque (red lines) with those that incorporate results from BLP18 (black solid lines) and the tracks obtained by implementing in our code the new recipes provided in Chrenko et al. (2024) (violet solid lines). For planets initially located beyond the water ice-line (i.e., beyond 3 au), Figure 11 illustrates that, although the overall formation tracks remain similar, the dust torque appears more effective at larger distances when employing the prescriptions of Chrenko et al. (2024). This results in a slightly more pronounced delay of rapid inward migration for those planets. In contrast, for planets that begin their formation inside the water ice-line, the effects of the dust torque become markedly more significant. This is because, according to the results in Chrenko et al. (2024) the positive magnitude of the dust torque is notably larger for low-mass planets and small Stokes numbers when pebble accretion is taken into account. Consequently, these planets migrate outward significantly during the early stages of formation. In addition, in this case all the inner planets cross the water-ice line and their formation is influenced by the accretion of ice-rich pebbles with larger Stokes numbers (as in the case of the metallic disk). Eventually, at some point along their formation pathways, the gas torque begins to dominate, leading the planets to migrate inward. The fact that the dust torque is more efficient for inner planets aligns well with the estimates provided by Chrenko et al. (2024), who employed a uniform pebble flux (see their Figure 7).
@@
Figure 11: Planet formation tracks for the fiducial disk –as in Figure 2– including in this case the impact of pebble accretion onto the dust torque (violet lines).
@@ 5 Summary and Discussion
The studies by BLP18, Regály (2020), and, more recently by Chrenko et al. (2024), have demonstrated that the asymmetric distribution of pebbles surrounding a planet can exert a significant torque on it, even though solid materials make up only a small fraction of the protoplanetary disk’s mass. Building on these findings (as well as on our previous work, Guilera et al., 2023), in this study we have quantified for the first time the influence of dust torque on the migration of planets growing through pebble accretion until either the disk dissipates or, conservatively, the planets reach a mass of $10~{}M_{\oplus}$. We accomplished this goal with PLANETALP, a comprehensive framework for modeling the formation history of planets that incorporates dust growth and evolution (Ronco et al., 2017; Guilera et al., 2017, 2019; Venturini et al., 2020b; Guilera et al., 2021).
We first employed a fiducial disk model with a moderate mass of $0.05M_{\odot}$ and a low $\alpha$-viscosity parameter of $10^{-4}$. Previous works on dust growth and evolution suggest that the efficiency of planet formation via pebble accretion is highly sensitive to the level of turbulence in the disk (e.g., Venturini et al., 2020b, a; Drazkowska et al., 2021; Guilera et al., 2021). On observational grounds, ALMA data suggest low $\alpha$-viscosity values (e.g., Dullemond et al., 2018; Flaherty et al., 2020). Therefore, based on these results and the findings of Venturini et al. (2020a) and Guilera et al. (2021), we considered a low $\alpha$-viscosity value, which enhances planetary formation efficiency.
@@
Impact of Dust Torque in Massive and Metal Rich Disks – We also investigated the role of the dust torque in the migration of planets within disks characterized by higher masses and metallicities. In both scenarios, we found that the mass-weighted mean pebble Stokes numbers across the disk remained comparable to those in the fiducial case. This similarity arises from the fact that dust growth is limited by fragmentation in the inner disk and by radial drift in the outer regions, resulting in minimal differences between these models. Consequently, the outcomes for these scenarios are qualitatively similar to the fiducial case. Although the dust torque plays a crucial role in the formation of planets initially located inside the water ice-line, it is less significant for those that start to form beyond the ice-line. This is due to the increased dust-to-gas mass ratio inside the water ice-line and the corresponding decrease of it outside, as a consequence of dust radial drift.
Notably, the disk with an initial metallicity of $Z=0.02$ exhibits a higher dust-to-gas ratio, resulting in a more pronounced impact of the dust torque. In this case, planets starting at 0.5 au and 1 au migrate outward until they reach approximately $2~{}M_{\oplus}$, while the planet initially at 2 au crosses the water ice-line. This change not only affects the final mass of the planet but also alters its composition. A similar trend was observed in Guilera et al. (2021) when updated prescriptions for the thermal torque were implemented.
@@
For planets initiating their formation within the water ice line, the influence of dust torque becomes more pronounced compared to the findings from BLP18. This is primarily because, as Chrenko et al. (2024) demonstrated, the dust torque tends to increase for low-mass planets with low Stokes numbers when pebble accretion is taken into account.
For planets initially positioned beyond the water ice line, we observe that considering the effects of pebble accretion on the dust torque has minimal influence on the formation tracks. This is even though the dust torque incorporating accretion is larger in magnitude than the one calculated using BLP18’s results. This occurs primarily because of the brief timescales for planet formation and the rapid decrease in the dust-to-gas mass ratio beyond the water-ice line. Furthermore, incorporating solid accretion enhances the magnitude of the dust torque at larger distances. Moreover, we note that the impact of the dust torque, when including accretion, seems to increase at larger distances from the star. This enhancement might stem from the different dependence of the dust torque on the normalized torque parameters, contrasting our approach (using the results from BLP18) with those from Chrenko et al. (2024, see their discussion in Sec. 3.5).
@@
We also note that the planets we modeled start forming at the beginning of the simulations, specifically at $t=0$ yr. As demonstrated by Drazkowska et al. (2021) and Guilera et al. (2021), if planetary embryos do not start forming in the early phases of disk evolution, pure pebble accretion becomes inefficient due to diminishing pebble flux over time caused by dust radial drift. Voelkel et al. (2022) incorporated a model for planetesimal formation through pebble flux and embryo formation from planetesimal growth in a framework similar to the one we use in this study. Their findings revealed that multiple generations of embryos can form prior to disk dissipation, particularly within the water ice-line, and that earlier generations primarily grow via pebble accretion, while later ones tend to grow through planetesimal accretion. More recently, Kaufmann et al. (2025) studied the formation of first embryos via planetesimal and pebble accretion in planetesimal rings forming by the streaming instability mechanisims. The key findings indicate that although the formation of planetary embryos from filaments can occur within the disk’s lifetime, it is highly sensitive to the distance from the central star. In the inner regions of the disk, planetesimal rings generated by the streaming instability can rapidly give rise to protoplanetary embryos, whereas in the outer disk, embryo formation requires substantially more time. We plan to incorporate these phenomena in future studies to examine further the impact of the dust torque on different generations of formed planets.
Finally, as we mentioned in our previous work (Guilera et al., 2023), our results inherently carry the same limitations as those present in the torque computations of BLP18. We note that the numerical simulations employed in BLP18 are 2D, and thus do not account for the vertical structure of the disk. We also note that the back-reaction of dust on the gas (i.e., feedback) is neglected in BLP18 (also in Chrenko et al., 2024). This assumes that the solid-to-gas density ratio remains low both in the bulk of the disk and in the vicinity of the planet. Nevertheless, Regály et al. (2025) showed recently that the inclusion of the back-reaction of solid material tends to increase the dust torque, especially for very low-mass planets (of the order of Mars mass).
Another simplification in BLP18 is the omission of planetary migration, which alters the relative motion between solids and the planet and may significantly affect the asymmetry in the dust distribution near the planet. This, in turn, could either enhance or reduce the resultant torque. If migration induces eccentricity in the planet’s orbit, the torque behavior may change further, underscoring the need for additional hydrodynamical studies to explore these effects.
@@ 6 Conclusions
Planets initially inside the water ice-line – In this case, the dust torque represents a significant contribution to the total torque due to the increased dust-to-gas mass ratio resulting from dust drift from the outer disk. In fact, at early stages, the dust torque can dominate the total torque acting on the planet, particularly in disks with initial metallicities slightly above the standard 1% ($\epsilon_{0}\gtrsim 0.02$). However, to achieve more accurate results, further detailed simulations extending to lower planet masses and Stokes numbers are necessary, especially when considering the recent results from Chrenko et al. (2024). For these planets, we identify three primary pathways:
@@
iii) when applying the prescriptions from Chrenko et al. (2024), all cases show that the planets achieve larger masses. In this scenario, the dust torque plays a more significant role in pushing (torquing) the planets outward, enabling them to cross the water-ice line and facilitating their formation through the accretion of larger pebbles, which also alters the planets’ composition by incorporating ice-rich pebbles.
@@
In conclusion, based on our results, we assert that the dust torque should not be ignored in models of planet formation, especially in research centered on the formation of terrestrial-like planets through pebble accretion. In this regard, the dust torque can significantly influence the overall torque experienced by planets undergoing pebble accretion, thereby impacting their migration patterns and, ultimately, their formation timescale, mass, and composition.
@@ 7 Acknowledgments
We thank the referee for a useful report with suggestions that helped us improve this manuscript. OMG and M3B are partially supported by PIP 2971 from CONICET (Argentina) and by PICT 2020-03316 from ANPCyT (Argentina). OMG and M3B also thank Juan Ignacio Rodriguez from IALP for the computation managing resources of the Grupo de Astrofísica Planetaria de La Plata. PBL acknowledges support from FONDECYT project 1231205. The research leading to this work was supported by the Independent Research Fund Denmark via grant ID 10.46540/3103-00205B attributed to MP.
@@ Appendix A Dust torque from the dust distribution
In this appendix, we compute the dust torque from the dust distribution instead of using only the mass-weighted mean Stokes number of such distribution. Even though it is a more expensive computational approach, it is useful to analyze if there is a significant difference with respect to the simpler approach of considering a mass-weighted mean Stokes number to characterize the dust distribution. To achieve this, we calculate the dust torque at the planet’s location for 200 dust species that are evenly spaced logarithmically between $\text{r}_{\text{d}}^{\text{min}}=1~{}\mu$m ($\text{St}_{\text{min}}$) and $\text{r}_{\text{d}}^{\text{max}}$ ($\text{St}_{\text{max}}$) (see Guilera et al., 2020, for details). We note again that in our model – which is based on the works of Birnstiel et al. (2012) and Drazkowska et al. (2016) – most of the mass remains in larger particles. We compute the dust-to-gas ratio corresponding to the i-species given by
@@
In Figure 12, we show a comparison of the planet formation tracks for the case where the dust torque is computed using the mass-weighted mean Stokes number of the dust distribution (the black lines) and the case where the dust torque is calculated for each dust species and then added up to compute the total dust torque (the multi-species case represented by the blue lines). We also plot the case where the dust torque is not considered in the total torque acting on the planets (the red lines). For the planets initially located inside the water ice-line, the planet formation tracks for the cases when the dust torque is considered are identical. This is due to the fact that inside the water ice-line the maximum Stokes number $\text{St}^{\text{max}}$ of the dust distribution is always lower than 0.01 (see the top panel of Figure 8). This also implies that the mass-weighted mean Stokes number is also lower than 0.01. Since we are not extrapolating outside the values of the dust torque map computed in BLP18, for St lower than 0.01 we are using this value to compute the dust torque. Hence, in the multi-species approach the dust torque for each dust species is the same, and the total dust torque is identical to the dust torque computed using the mass-weighted mean Stokes number. On the other hand, beyond the water ice-line the maximum Stokes number $\text{St}^{\text{max}}$ of the dust distribution reaches values larger than 0.01. Thus, in the multi-species approach not all the species contribute with the same value of the dust torque (particles with larger St contribute with larger dust torques). For those planets initially located beyond the water ice-line, we can see that there is a slight difference between the planet formation tracks when the dust torque is considered. Using the mass-weighted mean Stokes number of the dust distribution to calculate the dust torque slightly overestimates its effect compared to calculating the total dust torque based on the entire dust distribution. Nonetheless, in all scenarios, the dust torque contributes to delaying inward planet migration relative to cases where the dust torque is not taken into account.
Figure 12: Comparison of the planet formation tracks for the case of an initial disk metallicity of 0.02. The red curves correspond to the case where the total torque is given solely by the gaseous torque (where we use the recipes from Jiménez & Masset, 2017). The black and blue lines correspond to the case where the dust torque is considered. The black lines illustrate the scenario in which the dust torque is calculated using the mass-weighted mean Stokes number of the dust distribution, whereas the blue lines depict the scenario where the total dust torque is determined by summing the individual torque contributions from the 200 species within the distribution.